# Large language models surpass human experts in predicting neuroscience results

### 04 Mar 2024 v1

@@
Xiaoliang Luo Akilles Rechardt Guangzhi Sun Kevin K. Nejad Felipe Yáñez Bati Yilmaz Kangjoo Lee Alexandra O. Cohen Valentina Borghesani Anton Pashkov Daniele Marinazzo Jonathan Nicholas Alessandro Salatiello Ilia Sucholutsky Pasquale Minervini Sepehr Razavi Roberta Rocca Elkhan Yusifov Tereza Okalova Nianlong Gu Martin Ferianc Mikail Khona Kaustubh R. Patil Pui-Shee Lee Rui Mata Nicholas E. Myers Jennifer K Bizley Sebastian Musslick Isil Poyraz Bilgin Guiomar Niso Justin M. Ales Michael Gaebler N Apurva Ratan Murty Chloe M. Hall Jessica Dafflon Sherry Dongqi Bao Bradley C. Love
@@ Acknowledgments
This work was supported the ESRC (ES/W007347/1), Microsoft (Accelerate Foundation Models Research Program), and a Royal Society Wolfson Fellowship (18302) to B.C.L. We thank Mona Garvert, Pradeep Reddy Raamana, Todd Hare, Yoav Kessler, and Oliver Robinson for their assistance.
@@
One concern regarding LLMs outperforming human experts on BrainBench is the possibility that LLMs were exposed to the original abstracts during their pre-training. If LLMs have simply memorized the training data, they would naturally assign lower perplexity scores to the correct abstracts. For some models, like Galatica, their training set is open and we know they were not trained on BrainBench items.
@@
Bati Yilmaz, Anton Pashkov, Isil Poyraz Bilgin
@@
N Apurva Ratan Murty, Elkhan Yusifov, Kangjoo Lee
Valentina Borghesani, Sepehr Razavi, Kaustubh R. Patil
Michael Gaebler, Justin M. Ales, Guiomar Niso
Rui Mata, Mikail Khona, Roberta Rocca
Kevin K. Nejad, Alessandro Salatiello, Jonathan Nicholas
Daniele Marinazzo, Chloe M. Hall, Pui-Shee Lee
Sebastian Musslick, Nicholas E. Myers, Jennifer K Bizley
Sherry Dongqi Bao, Nianlong Gu
Jessica Dafflon, Bradley C. Love
@@
N Apurva Ratan Murty, Rui Mata, Isil Poyraz Bilgin
Tereza Okalova, Guiomar Niso, Ilia Sucholutsky
@@
Elkhan Yusifov, Jessica Dafflon, Felipe Yáñez
Jonathan Nicholas, Roberta Rocca, Valentina Borghesani
Sherry Dongqi Bao, Alexandra O. Cohen
Alessandro Salatiello, Xiaoliang Luo, Bradley C. Love

### 14 Mar 2024 v2

@@
Xiaoliang Luo Akilles Rechardt Guangzhi Sun Kevin K. Nejad Felipe Yáñez Bati Yilmaz Kangjoo Lee Alexandra O. Cohen Valentina Borghesani Anton Pashkov Daniele Marinazzo Jonathan Nicholas Alessandro Salatiello Ilia Sucholutsky Pasquale Minervini Sepehr Razavi Roberta Rocca Elkhan Yusifov Tereza Okalova Nianlong Gu Martin Ferianc Mikail Khona Kaustubh R. Patil Pui-Shee Lee Rui Mata Nicholas E. Myers Jennifer K Bizley Sebastian Musslick Isil Poyraz Bilgin Guiomar Niso Justin M. Ales Michael Gaebler N Apurva Ratan Murty Leyla Loued-Khenissi Anna Behler Chloe M. Hall Jessica Dafflon Sherry Dongqi Bao Bradley C. Love
@@ Results
As a final check (Fig. S.6; Materials and Methods), we confirmed that LLMs do not perform better on items published earlier in 2023 (e.g., Jan-2023 vs. Oct-2023), which addresses the concern that early items are more likely to have a preprint or other precursor appear in the training set that affected BrainBench performance. All our checks indicated that BrainBench items were novel for the LLMs.
@@ Acknowledgments
This work was supported the ESRC (ES/W007347/1), Microsoft (Accelerate Foundation Models Research Program), and a Royal Society Wolfson Fellowship (18302) to B.C.L. We thank Mona Garvert, Pradeep Reddy Raamana, Todd Hare, Yoav Kessler, and Oliver Robinson for their assistance. We thank the participants of the online study (see Table S.5).
@@
One concern regarding LLMs outperforming human experts on BrainBench is the possibility that LLMs were exposed to the original abstracts during their pre-training. If LLMs have simply memorized the training data, they would naturally assign lower perplexity scores to the correct abstracts.
@@
Finally, we analyzed the Spearman correlation between the publication dates of the abstracts that make up BrainBench test cases against the test cases’ difficulties to LLMs. This was to address the concern that early items are more likely to have a preprint or other precursor appear in the training set memorized by LLMs. If there was memorization, we would expect a negative correlation between publication date and difficulty. We determined difficulty by using the difference in perplexity between incorrect and correct abstracts for each test case.
@@
Bati Yilmaz, Isil Poyraz Bilgin, Anton Pashkov
@@
Elkhan Yusifov, N Apurva Ratan Murty, Kangjoo Lee
Valentina Borghesani, Sepehr Razavi, Justin M. Ales
Rui Mata, Michael Gaebler, Guiomar Niso
Leyla Loued-Khenissi, Anna Behler, Kaustubh R. Patil
Mikail Khona, Roberta Rocca, Kevin K. Nejad
Alessandro Salatiello, Jonathan Nicholas
Daniele Marinazzo, Chloe M. Hall
Pui-Shee Lee, Sebastian Musslick
Nicholas E. Myers, Jennifer K Bizley, Sherry Dongqi Bao
Nianlong Gu, Jessica Dafflon, Bradley C. Love
@@
N Apurva Ratan Murty, Ilia Sucholutsky, Guiomar Niso
Isil Poyraz Bilgin, Rui Mata, Tereza Okalova
@@
Elkhan Yusifov, Leyla Loued-Khenissi, Jessica Dafflon
Jonathan Nicholas, Felipe Yáñez, Roberta Rocca
Valentina Borghesani, Sherry Dongqi Bao
Alexandra O. Cohen, Alessandro Salatiello
Xiaoliang Luo, Bradley C. Love
@@
Gaia Molinaro, Judy Zhu, Majd Abdallah, Yuri G. Pavlov, Jong-eun Lee, Adam Harris, Zhaoning Li, Roman Kessler, Lexi Zhang, Maciej Szul, Pranjul Gupta, Sunreeta Bhattacharya, Jellina Prinsen, Celine Gallagher, Michael Anes, Maarten Laroy, Tobias Ackels, Carina Forster, Pedro Gonçalves, Tommy Mcconnell, Diane Whitmer, Debottam Kundu, Benjamin Pasquereau, Jeremy Manning, Maciej Szul, Ahmed Hussain, Nicolas Clairis, Ignacio Vega-Vásquez, Kun Chen, Jeremy Hogeveen, Sina Salehi, Suseendrakumar Duraivel , Edgar Guevara, Ziyao Zhang, Thomas J. Younts, Marek Muszyński, Leonardo Dalla Porta, Todd Gureckis, Parnian Rafei, Feng-Chun Chou, Keith Temple, Alp Altunkaya, Andrew Tan, Jin Ho Yun, Arnau Marin-Llobet, Brian Lord, Daniel Lindh, Simon Besson-Girard, Eren Irmak, Emin Çelik, Aman Maharjan, Irene Sophia Plank
Table S.5: Participants acknowledgment. List of participants.